News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 2 May 2003 9.00am EDT

During the last week the observing schedule was replanned to include an observation of the galactic black hole H1743-322 that was accepted as a Target Of Opportunity (TOO) on 25 April. The observing schedule was interrupted a second time at 7:36pm on 30 April due to a high radiation event that activated the science instrument safing sequence (SCS 107). The high radiation is thought to be associated with the currently active coronal stream as Chandra approached the radiation belts, as was the case on April 17. The loads were restarted at 5:15pm on May 1 and included the TOO. Observations of 3C296, B3 1659+399 and HD 163296 were impacted by these events (total 53Ks science time) and will be rescheduled in a later load.

The IRU-1 gyro-1 current remained at ~150mA (its yellow limit) with continued noisy behavior.

Real-time procedures were completed on 29 April to re-enable dither and re-enable SCS 29 and on May 1 to re-enable SCS 107 and disable the Momentum Unload subfunction. The April 29 actions were corrective and procedures will be assessed and adjusted accordingly.

The observing schedule for next week is shown below and includes an additional pad-time prior to radiation belt entry to mitigate the present elevated radiation conditions.

------------------------------------------
RCW103                 ACIS-I       May 4
SDSS J222845-075755    ACIS-S
DOAR 21                ACIS-S/HETG  May 5
Radiation Belts                     May 6
PDS 456                ACIS-S/HETG
SCO X-1                ACIS-S/HETG  May 8
Radiation Belts
MXB 1659-298           ACIS-S       May 9
G11.2-0.3              ACIS-S
MRK 279                HRC-S/LETG   May 10
HD110073               ACIS-I
NGC 1800               ACIS-S

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports